a. If Canace Company, with a break-even point at $329,400 of sales, has actual sales of $540,000, what is the margin of safety expressed (1) in dollars and (2) as a percentage of sales? Round the percentage to the nearest whole number.

1. $

2.   %

b. If the margin of safety for Canace Company was 40%, fixed costs were $1,725,600, and variable costs were 60% of sales, what was the amount of actual sales (dollars)?
(Hint: Determine the break-even in sales dollars first.)
$
Answer #1

a)

1.

Margin of safety = Actual sales - Break even sales

= 540,000 - 329,400

= $210,600

2.

Margin of safety (%) = Margin of safety/Actual sales

= 210,600/540,000

= 39%

b)

Margin of safety for Canace Company was 40%, fixed costs were $1,725,600, and variable costs were 60% of sales.

Since variable costs are 60% of sales, hence Contribution margin ratio must be 40%

Break even point ($) = Fixed cost/Contribution margin ratio

= 1,725,600/40%

= $4,314,000

Since margin of safety is 40%, hence break even must occur at 60% of sales.

Hence, break even point = Actual sales x 60%

4,314,000 = Actual sales x 60%

Actual sales = 4,314,000 x 100/60

= $7,190,000

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubt. Thanks